Moody Corporation uses a job-order costing system with a plantwide predetermined overhead rate based on machine-hours. At theRequired 1 Required 2 Required 3 Required 4 Compute the plantwide predetermined overhead rate. (Round your answer to 2 decimaRequired 1 Required 2 Required 3 Required 4 Compute the total manufacturing cost assigned to Job 400. (Round your intermediatRequired 1 Required a Required 2 Required 2 Required 3 Required 3 Required 4 required a If Job 400 includes 60 units, what isRequired 1 Required 2 Required 3 Required 4 If Moody uses a markup percentage of 110% of its total manufacturing cost, then w
Answer #1

1.

Predetermined overhead rate = Variable overhead cost per machine hour + Fixed overhead cost/Estimated machine hours

= 4.30 + 651,000/151,000

= 4.30 + 4.31

= $8.61 per machine hour

Predetermined overhead rate $8.61 per MH

2.

Overhead applied = Actual machine hours used x  Predetermined overhead rate

= 38 x 8.61

= $327.18

Job cost sheet

Direct material

390

Direct labor

250

Overhead applied

327.18

Total cost

$967.18

Total manufacturing cost $967.18

3.

Unit product cost = Total manufacturing cost / Number of units

= 967.18/60

= $16

Unit product cost $16

4.

Markup percentage = 110% of total job cost

Mark up = 967.18 x 110%

= $1,063.90

Total sales value = Total cost + Mark up

= 967.18+1,063.90

= $2,031.08

Selling price per unit  = Total sales value /Number of units

=2,031.08/60

= $34

Selling price per unit $34

Note: Exact answer may slightly differ due to rounding off

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks